OpenStudy (anonymous):

(-1, 3) and perpendicular to y = -one thirdx + 7. y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (texaschic101):

y = -1/3x + 7 do you know what the slope is in this equation ?

OpenStudy (texaschic101):

y = mx + b...the slope is in the m position. Therefore, the slope of this equation is -1/3. However, we are looking for a line perpendicular, so we need to find the negative reciprocal of the slope in the equation. All that means is you " flip " the slope and change the sign. So, the slope we need is the negative reciprocal of -1/3, which is 3/1 or just 3. Understand so far ?

OpenStudy (texaschic101):

now we will use point slope form...y - y1 = m(x - x1) slope(m) = 3 (-1,3) x1 = -1 and y1 = 3 now lets sub our info into the formula y - 3 = 3(x - (-1) y - 3 = 3(x + 1) y - 3 = 3x + 3 y = 3x + 3 + 3 y = 3x + 6 <-- your answer any questions just ask :)